OpenStudy (anonymous):

5- 10 l0.8 - 3/5l A. -5 B. -1 C. 3 D. 5

OpenStudy (anonymous):

3/5 = 0.6 0.8 - 0.6 = 0.2 10* 0.2 = 2 5 - 2 = 3 3 is your answer